OpenStudy (anonymous):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically. (4 points)

OpenStudy (anonymous):

Let's name "x" the number of years. we know that 15 +15*1.20^{x} =40+x let's try after 5 years: Roger would have 45 cars and Cory 52 ( i just guessed 5, and then I calculated 15 +15*1.20^{5} ): so, no, it has to be less than 5. let's try 4: Roger will have 44, and cory 15+31=45, so that's quite good asnwer! Let's try with 3: Roger will have 43, and Cory 15+ 25=40, so 4 years are better: after approximately 4 years. (perhaps a better strategy would be to just calculate after each year starting at 1 and going up...)

OpenStudy (anonymous):

I need to know how 4 is the correct answer?

OpenStudy (melissa_something):

He just explained, if anything plug 4 into the original and see if it works

OpenStudy (anonymous):

It doesn't